OpenStudy (junyang96):

Using the solar constant(approx 1.361kW/m^2) estimate the rate at which the whole Earth receives energy from the Sun

OpenStudy (anonymous):

Earth is a sphere so that its surface area = 4pi r^2. and Sun will hit only half its area at a time so cross sectional area =2pi r^2. rate = E/A thus Energy = rate *A

OpenStudy (junyang96):

but the answer given was 1.73e17 which is he result of solar constant*pi r^2 similar question here https://answers.yahoo.com/question/index?qid=20091108092559AAbV0tM

OpenStudy (anonymous):

Yea that is right as rate should be parallel to the direction of area ,so it would be the cross sectional area which is pi r^2 or 2pi r^2 (dot product with the rate).

OpenStudy (anonymous):